OpenStudy (anonymous):

How do you write the equation for this: A bus traveled on a level road for 3 hours at an average spped 20 miles per hour faster than it traveled on a winding road. The time spent on the winding road was 4 hours. Find the average speed on the level road if the entire trip was 305 miles.

OpenStudy (anonymous):

yikes i need pencil and paper cannot do this in my head today

OpenStudy (anonymous):

I think I can solve it I'm just not sure how to write the equation in the d=rt form.

OpenStudy (anonymous):

x = speed on the winding road \[305 = 3(20 + x) + 4x = 60 + 7x\]\[x = 35\]Speed on level road is \[35 + 20 = 55\]

OpenStudy (anonymous):

that was snappy

OpenStudy (anonymous):

Thank you very much!!!!!